OpenStudy (tiffanymak1996):

find the area of the region enclosed by the rose r=6cos3t

OpenStudy (lgbasallote):

dont you use integrals?

OpenStudy (ganpat):

ohhh, Rose curve.. right ?? compare it with r =acos(Kt) Area = (Pi * a2) / 2 .............................. If K is even,, Area = (Pi * a2) / 4 .............................. if K is odd.. i guess.. :-p

OpenStudy (ganpat):

@lgbasallote : what say ?? its been long...

OpenStudy (tiffanymak1996):

i don't get it

OpenStudy (anonymous):

integrate within the limits of t....

OpenStudy (ganpat):

r=6cos3t comparing it with.. r =acos(Kt) a = 6 k = 3.. K is odd number.. SO, Area = (Pi * a2) / 4 = (Pi * 6 sq) / 4 = (3.14 * 36) / 4 = 28.26... ????????????????????

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Rose_(mathematics)

OpenStudy (ganpat):

oh.. m right :)..

OpenStudy (anonymous):

you want solve it by integration?

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

t is angle ?

OpenStudy (tiffanymak1996):

yes

OpenStudy (tiffanymak1996):

first: graph the curve: http://www.wolframalpha.com/input/?i=r%3D6cos3t+graph+rose

OpenStudy (tiffanymak1996):

since the area enclosed by the rose has 6 equal parts, we'll find one and multiply it by 6.

OpenStudy (tiffanymak1996):

using the formula for polar curve's area: \[A=\int\limits_{b}^{a} \frac{ 1 }{ 2 }r^2 dt\]

OpenStudy (tiffanymak1996):

so, \[Area = 6 \int\limits_{0}^{\frac{ \pi }{ 3 }} \frac{ 1 }{ 2 }(6cost)^2 dt\]

OpenStudy (tiffanymak1996):

\[= 108\int\limits_{0}^{\frac{ \pi }{ 3 }} \cos^2 3t dt\]

OpenStudy (tiffanymak1996):

\[= 36 \int\limits_{0}^{\frac{ \pi }{ 3 } } \cos^23t d(3t)\]

OpenStudy (tiffanymak1996):

\[=36[\frac{ 1 }{ 2 } (3t) + \frac{ 1 }{ 4 } \sin (6t)]_{0}^{\frac{ \pi }{ 3 }}\]

OpenStudy (tiffanymak1996):

\[= 36[ (\frac{ \pi }{ 2 }+0)-(0+0)]\]

OpenStudy (tiffanymak1996):

\[=18 \pi\]

OpenStudy (tiffanymak1996):

and there you go, all done!

OpenStudy (anonymous):

This stunning

OpenStudy (anonymous):

You are wonderful , but there are something i want ask about can i ?

OpenStudy (tiffanymak1996):

go on ahead!

OpenStudy (anonymous):

first how do you know the shape of the graph ?

OpenStudy (anonymous):

you dont know any mathematical way to know it ?!

OpenStudy (tiffanymak1996):

the question said rose, so it's a flower like curve, for how many petals you know it by r=a cos (nt) if n=3 then 3 petals

OpenStudy (tiffanymak1996):

if n=2 then 4 petals, n=3 3 petals, n=4 8 petals, n=5 5 petals, n=6 12 petals, n=7 7 petals and so on

OpenStudy (tiffanymak1996):

this sequence works only for a rose, for other polar curves, refer to your book, or other internet materials.

OpenStudy (anonymous):

thats perfect , Iam sorry ,I studied integration past year alone and did not study these rules

OpenStudy (tiffanymak1996):

nah, this isn't within the integration department, it's with the parametric and polar curves part, i think you learn this after integration.